# Wasserstein distances and divergences of order $p$ by quantum channels

### v1

@@ 1.1. Motivation and main result
It is a general phenomenon that concepts and notions that are well-established in the classical commutative world do not have a unique “best” extension in the non-commutative world but there are many possible ways of generalization with pros and cons. This is the case concerning optimal transportation as well. Non-commutative optimal transport is a flourishing research field these days with several essentially different promising approaches such as that of Biane and Voiculescu (free probability) [1], Carlen, Maas, Datta, Rouzé, and Wirth (dynamical theory) [11, 10, 9, 13, 12, 51], Caglioti, Golse, Mouhot, and Paul (quantum many-body problems) [6, 7, 25, 26, 31, 30, 27], De Palma and Trevisan (quantum channels) [15, 16], Życzkowski and his collaborators [23, 52, 53, 2], and Duvenhage [18, 20, 19]. Separable quantum Wasserstein distances have also been introduced recently [47]. For a comprehensive overview of a substantial part of the above mentioned current approaches to non-commutative optimal transport, the reader is advised to consult the book [39].
@@ 1.2. Basic notions, notation
where $\pi$ runs over all possible couplings of $\mu$ and $\nu.$ We say that $\pi\in\mathrm{Prob}(\mathbb{R}^{K}\times\mathbb{R}^{K})$ is a coupling of $\mu$ and $\nu$ (in notation: $\pi\in\mathcal{C}(\mu,\nu)$) if the marginals of $\pi$ are $\mu$ and $\nu,$ that is, $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}f(x)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}f(x)\mathrm{d}\mu(x)$ and $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}g(y)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}g(y)\mathrm{d}\nu(y)$ for all $f,g\in C_{b}(\mathbb{R}^{K}).$ A consequence of the tightness (that is, sequential compactness in the weak topology) of $\mathcal{C}(\mu,\nu)$ and the lower-semicontinuity of $c$ is that there is a coupling (in other words: transport plan) $\pi_{0}\in\mathcal{C}(\mu,\nu)$ that minimizes (1), see, e.g., [48, Thm. 4.1.].
@@ 2.1. A quantum mechanical optimal transport problem generalizing [15]
In transport theory, a usual “finite energy condition” ensuring that $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\pi(x,y)$ is finite for every coupling of $\mu$ and $\nu$ is that $c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu)$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu)$ — see, e.g., [48, Remark 4.5]. Accordingly, here we impose the following finite energy condition. Let $\mu_{k}$ and $\nu_{k}$ be the laws of $A_{k}$ in the states $\rho$ and $\omega$ respectively, that is, $\mathrm{d}\mu_{k}(x_{k})=\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]$ and $\mathrm{d}\nu_{k}(y_{k})=\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$ for every $k\in\{1,\dots,K\},$ and let $c(x,y)$ satisfy $0\leq c(x,y)\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu_{1}\otimes\dots\otimes\mu_{K})$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu_{1}\otimes\dots\otimes\nu_{K}).$ Under this assumption, $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]$ is finite for any $\pi\in\mathcal{C}(\rho,\omega).$ Indeed, in this case
@@
Then, according to the compactness result concerning quantum transport plans [15, Proposition 8], there is a subsequence $\left\{\Pi_{n_{j}}\right\}_{j\in\mathbb{N}}$ and an element $\widetilde{\Pi}\in\mathcal{C}(\rho,\omega)$ such that $\Pi_{n_{j}}\to\widetilde{\Pi}$ in trace norm. Consequently, $\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}$ converges to $\mathbb{P}_{\left(\widetilde{\Pi}\right)}^{(\mathcal{A})}$ set-wise. Indeed, for any Borel set $B\subset\mathbb{R}^{2K}$ the convergence
@@
on the set of Borel probabilities on $\mathbb{R}^{K}\times\mathbb{R}^{K}$ with respect to the set-wise convergence. This is a standard argument (see, e.g., [48, Lemma 4.3]) which is based on the lower semi-continuity of $(x,y)\mapsto c(x,y)$ itself. As $c$ is non-negative and lower semi-continuous, it is the pointwise limit of a monotone increasing family of non-negative, continuous and bounded functions $\left\{c_{r}\right\}_{r\in\mathbb{N}}$ on $\mathbb{R}^{K}.$ By Beppo Levi’s monotone convergence theorem,
@@
In the even more special case $p=q=2$ we get back the definition of the quadratic quantum Wasserstein distance introduced by De Palma and Trevisan in [15, Definition 8].
@@ 6. $p=2$ as a possible threshold for the triangle inequality
An important feature of the quadratic Wasserstein distances is that the distance of a state $\rho$ from itself (which may be positive) is always realized by the identity channel — see [15, Corollary 1]. Therefore, the quadratic Wasserstein distance of a state $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ from itself admits the explicit form
@@ 7. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 19
is a special case of the monotonicity version of Lieb’s concavity — see, e.g., [8, Theorem 1]. We can use (117) for a further estimation in the following way:

### v2

@@ 1.1. Motivation and main result
It is a general phenomenon that concepts and notions that are well-established in the classical commutative world do not have a unique “best” extension in the non-commutative world but there are many possible ways of generalization with pros and cons. This is the case concerning optimal transportation as well. Non-commutative optimal transport is a flourishing research field these days with several essentially different promising approaches such as that of Biane and Voiculescu (free probability) [2], Carlen, Maas, Datta, Rouzé, and Wirth (dynamical theory) [12, 11, 10, 14, 13, 52], Caglioti, Golse, Mouhot, and Paul (quantum many-body problems) [7, 8, 26, 27, 32, 31, 28], De Palma and Trevisan (quantum channels) [16, 17], Życzkowski and his collaborators [24, 53, 54, 3], and Duvenhage [19, 21, 20]. Separable quantum Wasserstein distances have also been introduced recently [48], and a novel definition of quantum transport plans has been proposed by Beatty and Stilck França in [1] where they introduced the corresponding $p$-Wasserstein distances as well. For a comprehensive overview of a substantial part of the above mentioned current approaches to non-commutative optimal transport, the reader is advised to consult the book [40].
@@ 1.2. Basic notions, notation
where $\pi$ runs over all possible couplings of $\mu$ and $\nu.$ We say that $\pi\in\mathrm{Prob}(\mathbb{R}^{K}\times\mathbb{R}^{K})$ is a coupling of $\mu$ and $\nu$ (in notation: $\pi\in\mathcal{C}(\mu,\nu)$) if the marginals of $\pi$ are $\mu$ and $\nu,$ that is, $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}f(x)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}f(x)\mathrm{d}\mu(x)$ and $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}g(y)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}g(y)\mathrm{d}\nu(y)$ for all $f,g\in C_{b}(\mathbb{R}^{K}).$ A consequence of the tightness (that is, sequential compactness in the weak topology) of $\mathcal{C}(\mu,\nu)$ and the lower-semicontinuity of $c$ is that there is a coupling (in other words: transport plan) $\pi_{0}\in\mathcal{C}(\mu,\nu)$ that minimizes (1), see, e.g., [49, Thm. 4.1.].
@@ 2.1. A quantum mechanical optimal transport problem generalizing [16]
In transport theory, a usual “finite energy condition” ensuring that $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\pi(x,y)$ is finite for every coupling of $\mu$ and $\nu$ is that $c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu)$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu)$ — see, e.g., [49, Remark 4.5]. Accordingly, here we impose the following finite energy condition. Let $\mu_{k}$ and $\nu_{k}$ be the laws of $A_{k}$ in the states $\rho$ and $\omega$ respectively, that is, $\mathrm{d}\mu_{k}(x_{k})=\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]$ and $\mathrm{d}\nu_{k}(y_{k})=\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$ for every $k\in\{1,\dots,K\},$ and let $c(x,y)$ satisfy $0\leq c(x,y)\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu_{1}\otimes\dots\otimes\mu_{K})$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu_{1}\otimes\dots\otimes\nu_{K}).$ Under this assumption, $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]$ is finite for any $\pi\in\mathcal{C}(\rho,\omega).$ Indeed, in this case
@@
Then, according to the compactness result concerning quantum transport plans [16, Proposition 8], there is a subsequence $\left\{\Pi_{n_{j}}\right\}_{j\in\mathbb{N}}$ and an element $\widetilde{\Pi}\in\mathcal{C}(\rho,\omega)$ such that $\Pi_{n_{j}}\to\widetilde{\Pi}$ in trace norm. Consequently, $\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}$ converges to $\mathbb{P}_{\left(\widetilde{\Pi}\right)}^{(\mathcal{A})}$ set-wise. Indeed, for any Borel set $B\subset\mathbb{R}^{2K}$ the convergence
@@
on the set of Borel probabilities on $\mathbb{R}^{K}\times\mathbb{R}^{K}$ with respect to the set-wise convergence. This is a standard argument (see, e.g., [49, Lemma 4.3]) which is based on the lower semi-continuity of $(x,y)\mapsto c(x,y)$ itself. As $c$ is non-negative and lower semi-continuous, it is the pointwise limit of a monotone increasing family of non-negative, continuous and bounded functions $\left\{c_{r}\right\}_{r\in\mathbb{N}}$ on $\mathbb{R}^{K}.$ By Beppo Levi’s monotone convergence theorem,
@@
In the even more special case $p=q=2$ we get back the definition of the quadratic quantum Wasserstein distance introduced by De Palma and Trevisan in [16, Definition 8].
@@ 6. $p=2$ as a possible threshold for the triangle inequality
An important feature of the quadratic Wasserstein distances is that the distance of a state $\rho$ from itself (which may be positive) is always realized by the identity channel — see [16, Corollary 1]. Therefore, the quadratic Wasserstein distance of a state $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ from itself admits the explicit form
@@ 7. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 19
is a special case of the monotonicity version of Lieb’s concavity — see, e.g., [9, Theorem 1]. We can use (117) for a further estimation in the following way: